How many liters of pure alcohol must be added to 20 liters of a 30% solution to obtain an 80% solution?
Answer by TimothyLamb(4379)

30(20) + 100v = 80(20 + v)
600 + 100v = 1600 + 80v
20v = 1000
v = 50
---
answer:
add 50 liters of pure alcohol
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php